OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 5 meters and legs that measure 3 meters and 4 meters. What is the surface area of this prism if the height of the prism is 6 meters? When I tried to do the problem, I got 102 squared and got it wrong. How do I figure it out correctly?

OpenStudy (anonymous):

surface area is the sum of the individual sides's areas

OpenStudy (anonymous):

so you start with the base

OpenStudy (anonymous):

4*3/2 = 6

OpenStudy (anonymous):

now do the vertical sides above the two legs

OpenStudy (anonymous):

4 * 6/2 = 12 3 * 6/2 = 9

OpenStudy (anonymous):

now do the face above the hypotenuse

OpenStudy (anonymous):

there are a few ways of doing this

OpenStudy (anonymous):

the easiest would probably be with trigonometric functions. Have you covered those yet?

OpenStudy (anonymous):

Yes. You mean sin and cosine, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we know the base is 5, the harder part is finding the height.

OpenStudy (anonymous):

actually, now that I think about it this method does not require trig anyway.

OpenStudy (anonymous):

but whatever

OpenStudy (anonymous):

xD Okay

OpenStudy (anonymous):

The problem also says "What is the surface area of this prism if the height of the prism is 6 meters? " if that is height we are trying to find.

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

we need to find the height of the face, not of the prism

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):